# CITRAS: Covariate-Informed Transformer for Time Series Forecasting

### v2

@@ Appendix D Computational Efficiency
Besides the theoretical analysis, we demonstrate the empirical model efficiency on the BS dataset and PEMS08 dataset. We use the best model configuration for each model and evaluate using the same batch size for a fair comparison. The results in the BS dataset underscore the efficient model design of CITRAS. In the PEMS08 dataset, we observe the trade-off between training time and model performance in terms of MSE. Among them, CITRAS achieves the best performance without significant compromise in efficiency.
@@
Appendix E Error Bar
We report the standard deviation of CITRAS performance on covariate-informed forecasting with future known covariates under five runs with different random seeds. We also refer to the result of some competitive baselines from Table 1. The result in Table 8 exhibits that the performance of CITRAS is stable.
Table 8: Error bar of CITRAS on the covariate-informed forecasting with future covariates. The results of baselines are from Table 1 for reference.
@@ Appendix E Error Bar
0.338
0.357
0.286
0.318
0.358
0.360
Appendix F Extra Results
F.1 Comparison with Pre-trained Models
We further compare CITRAS with large pre-trained models under covariate-informed settings with future covariates. For baselines, we choose Moirai [46] (moirai-1.1-R-large) and Moirai-MoE [27] (moirai-moe-1.0-R-base), which are state-of-the-art pre-trained models that natively support future covariates. For a fair comparison, we present the zero-shot results of these models and the full-shot results of Moirai, which is fine-tuned with the full training dataset for a maximum of five epochs with a learning rate $lr\in\{1e-5,1e-6,1e-7\}$. Table 9 shows that CITRAS consistently outperforms Moirai under full-shot settings. Note that the implementation of ChronosX [4], which fine-tunes univariate pre-trained models to integrate covariates, is not publicly available yet and therefore not comparable.
Setting Model
@@ F.1 Comparison with Pre-trained Models
F.2 Full Results of Multivariate Forecasting
The full results of multivariate forecasting in Section 4.2 are listed in Table 10. We adopt the rolling forecasting approach [29], where a model trained with predetermined input and forecasting lengths is evaluated with a longer forecasting length by iteratively integrating its output into the subsequent input. Following previous works [29, 30], we set the input length to 672, the training forecasting length to 96, the testing forecasting length to {96, 192, 336, 720}, and the patch length to 96.
Table 10: Full results of the multivariate forecasting, where the input length is set to 672, and the forecasting length is set to {96, 192, 336, 720}. The evaluation protocol is the same as in Table 2. Results are cited from Timer-XL [30] if available; otherwise reproduced.

### v3

@@ B.2 Computational Efficiency
Besides the theoretical analysis, we demonstrate the empirical model efficiency on the BS dataset and PEMS08 dataset in Fig. 6. We use the best model configuration for each model and evaluate using the same batch size for a fair comparison. The results in the BS dataset underscore the efficient model design of CITRAS. In the PEMS08 dataset, we observe the trade-off between training time and model performance in terms of MSE. Among them, CITRAS achieves the best performance without significant compromise in efficiency.
@@
Appendix C Extra Results
C.1 Error Bar
We report the standard deviation of CITRAS performance on covariate-informed forecasting with future known covariates under five runs with different random seeds. The result in Table 6 exhibits that the performance of CITRAS is stable.
@@ C.1 Error Bar
Table 6: Error bar of CITRAS on the covariate-informed forecasting with future covariates.
C.2 Comparison with Pre-trained Models
We further compare CITRAS with large pre-trained models under covariate-informed settings with future covariates. For baselines, we choose Moirai (Woo et al. 2024) (moirai-1.1-R-large) and Moirai-MoE (Liu et al. 2024a) (moirai-moe-1.0-R-base), which are state-of-the-art pre-trained models that natively support future covariates. For a fair comparison, we present the zero-shot results of these models and the full-shot results of Moirai, which is fine-tuned with the full training dataset for a maximum of five epochs with a learning rate $lr\in\{1e-5,1e-6,1e-7\}$. Table 7 shows that CITRAS consistently outperforms Moirai under full-shot settings. Note that the implementation of ChronosX (Arango et al. 2025), which fine-tunes univariate pre-trained models to integrate covariates, is not publicly available yet and therefore not comparable.
Setting
Model
@@ C.2 Comparison with Pre-trained Models
C.3 Full Results of Multivariate Forecasting
The full results of multivariate forecasting in Section 4.2 are listed in Table 8. We adopt the rolling forecasting approach (Liu et al. 2024c), where a model trained with predetermined input and forecasting lengths is evaluated with a longer forecasting length by iteratively integrating its output into the subsequent input. Following previous works (Liu et al. 2024c, 2025), we set the input length to 672, the training forecasting length to 96, the testing forecasting length to {96, 192, 336, 720}, and the patch length to 96.
@@ C.3 Full Results of Multivariate Forecasting
Table 8: Full results of the multivariate forecasting, where the input length is set to 672, and the forecasting length is set to {96, 192, 336, 720}. The evaluation protocol is the same as in Table 2. Results are cited from Timer-XL (Liu et al. 2025) if available; otherwise reproduced.
Appendix D Limitations and Future Work
In this paper, we explored effective ways to incorporate time-varying covariates into the forecasting process. We consider static covariates to be outside the scope of this paper and leave them for future work. Recently, numerous papers have tackled large-scale pre-training to achieve generalization performance in time series forecasting. Our model is based on a decoder-only Transformer trained through next-token prediction. This design is assumed to be suitable for large-scale pre-training, but its generalization capability has not yet been evaluated. In the future, we will explore the generalization potential of CITRAS in covariate-informed settings.